OpenStudy (anonymous):

Write an equation of the line normal to y= 1/(x+1) at x=2

OpenStudy (anonymous):

So first you need to find slope , dy/dx=-1/(x+1)^2=-1/9,This is slope of tangent to get slope of normal it is -1/dy/dx=sustituting two you get slope=9 so you also know a point (2,1/3) . Use slope point form to get the answer.

OpenStudy (anonymous):

wait how did u get 1/9

OpenStudy (anonymous):

err slope of tangent=-1/9 so slope of normal=-1/(slope of tangent)=-1/(-1/9)=9.

OpenStudy (anonymous):

Did you get what i have done?

OpenStudy (anonymous):

yes and i got y = 9x - (55/3)

OpenStudy (anonymous):

but how did u get taht point (2, 1/3)

OpenStudy (anonymous):

you want to find normal at x=2 no so the normal has to pass through the the curve at the point where x=2, now you know y=1/(x+1) put x=2 in this you get y=1/3 hence the point...

OpenStudy (anonymous):

ohhh ok XD....srry sheesh im soo slow

OpenStudy (anonymous):

thanks for your help

OpenStudy (anonymous):

No problem....